The contribution margin ratio is 25% for Crowne Company and the break-even point in sales is $220,000. If Crowne Company's target operating profit is $62,000, sales would have to be:

Multiple Choice

  • $282,000.

  • $276,000.

  • $468,000.

  • $248,000.
Answer #1

Breakeven=Fixed expenses/Contribution margin ratio

Hence fixed expenses=(220,000*0.25)=$55,000

Target contribution margin for target operating profit=Fixed expenses+Target profit

=(55000+62000)=$117,000

Hence target sales=117,000/0.25

=$468,000